OpenStudy (blehblehbleh):

Will medal and fan. Really stumped on this crap. Which ordered pairs are a solution to the equation? y = 3x - 4 Select all that apply. (4,2) (1, -1) (-1, -7) (2, 2)

OpenStudy (wolfman77):

Okay, so whatever x equals, it has to go in order with y So, if x was 4, would y be 2? Lets see: 3*2(x) = 6 - 4 = 2 So, yeah, it would be. It's A) (4,2) Hope this helps! =)